Grade 10Electrostatics

Four charges equal to -Qare placed four corners of a square and a charge q is at centre and whole system is in equillibrium .What is value of q?

Barsat Kher
9 Years agoGrade 10

1 Answer

Vikas TU
9 Years ago
We have the framework which comprise of 4 point charges +q set at 4 corners of a square . The four of charges will be in balance if add up to electric potential vitality is zero.Hence, = Potential Energy because of all the four sets in which Potential vitality framework charges put at the side of the sides of square+PE because of each one of those sets in which charges are set at finishes of the slanting of square +Potential Energy because of sets shaped by any charge at the corner and charge 
 
Potential Energy of System = 4×1/4πε0 ×q2/a +2×1/4πε0 × q2/a√2 +4×1/4πε0 × qQ/a/√2 
 
Potential Energy of System = 1/4πε0×4q/a [q+q/2√2+Q√2 ] 
 
The framework will be in balance if 
 
Potential Energy of System = 0 
 
1/4πε0 ×(4q/a) [q+ q/2√2+Q√2 ] = 0 
 
q+ q/2√2+Q√2 = 0 
 
Q√2 = −(q+q/2√2 ) 
 
Q = −(q/√2 + q/4) 
 
Q = −q((2√2+1 )/4) 
 
Q = −q/4 (2√2 +1) 
 
We have the framework which comprise of 4 point charges +q put at 4 corners of a square . The four of charges will be in balance if add up to electric potential vitality is zero.Hence, = Potential Energy because of all the four sets in which Potential vitality framework charges set at the edge of the sides of square+PE because of every one of those sets in which charges are set at finishes of the askew of square +Potential Energy because of sets shaped by any charge at the corner and charge 
 
Potential Energy of System = 4×1/4πε0 ×q2/a +2×1/4πε0 × q2/a√2 +4×1/4πε0 × qQ/a/√2 
 
Potential Energy of System = 1/4πε0×4q/a [q+q/2√2+Q√2 ] 
 
The framework will be in harmony if 
 
Potential Energy of System = 0 
 
1/4πε0 ×(4q/a) [q+ q/2√2+Q√2 ] = 0 
 
q+ q/2√2+Q√2 = 0 
 
Q√2 = −(q+q/2√2 ) 
 
Q = −(q/√2 + q/4) 
 
Q = −q((2√2+1 )/4) 
 
Q = −q/4 (2√2 +1)